Project

General

Profile

Actions

20180302013 » History » Revision 3

« Previous | Revision 3/9 (diff) | Next »
Whyte, Colin, 13 March 2018 12:56


20180302013

Location : MICE meeting room R1 2nd Floor East at 14:00 pm

Phone access : phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • February jobs:
    • Tracker:
      • Cryostats have been disconnected. 1,2&3 ready to move. 4 trapped by vac lines. 3 Sumitomo systems stripped out
      • Helium window measurement - removed complete - move to tracker lab for metrology.
    • Diffuser extracted (survey complete) -- and moved to tracker lab. Need to verify thickness and composition of the diffuser. (TM, CMcW)
    • Target Controls test #1954 (Henry/Paolo 21st) Requires PC move into Tracker rack.
    • D1 field.
      • Limited plan from HN.
    • https://micewww.pp.rl.ac.uk/attachments/download/9660/2017-12-06_decommissioning-measurements_v3.pdf
      • Document stands; no updates.
  • March jobs:
    • SSU, SSD Magnet field map. 19th to 23rd March
      • equipment delivered to R9 - move to MICE hall later in week.
      • Cern team flights confirmed.
      • first survey in place afternoon Mon 19th - TM advised.
      • access passes for the CERN staff with Rose Hayes- complete?
    • QPS test today
    • SSU vacuum - north side turbo on - good vacuum?.
    • hall search.
  • collaborator equipment return to power - in progress - new crates required- JG
  • Next meeting - 20th Mar18.
  • R9 empty by 19th.
  • AOB

Updated by Whyte, Colin over 4 years ago · 3 revisions